# SubGrapher: Visual Fingerprinting of Chemical Structures

### v1

@@ 7.1.2 State-of-the-art Comparison
Figure 4: Qualitative evaluation. Examples of predicted functional groups are shown for images from patent documents (JPO, USPTO-10K-L, USPTO-Markush) and a scientific journal (JOC).
@@
On the USPTO-Markush dataset, SubGrapher again outperforms other methods in substructure F1-score. This superior performance results from the limitations of competing models: OSRA and MolGrapher are trained exclusively on molecular images, whereas MolScribe and DECIMER handle Markush structures but support only limited Markush features, specifically variable groups represented as abbreviations. In contrast, SubGrapher effectively manages complex Markush structures by focusing on relevant regions of the image and excluding Markush-related annotations.
@@ 7.2 Qualitative Evaluation
Figure 4 showcases examples of predicted molecules for images from various benchmark datasets. SubGrapher accurately recognizes functional groups in molecule images that contain captions or are of lower quality (Figure 4, row 1). It also recognizes functional groups in complex Markush structures (Figure 4, row 2) or unconventional drawings displayed in scientific publications (Figure 4, row 3). Additionally, our method maintain strong performances on extremely large molecules (Figure 4, row 4). Unlike image captioning methods such as DECIMER, our predictions preserve the spatial arrangement of substructures from the input image. This conveys useful information for human interpretation.
@@ 7.3 Visual Fingerprinting Evaluation
In this section, we evaluate SubGrapher and state-of-the-art methods for the retrieval of a molecule in a collection of molecule images.
Figure 5: Visual fingerprinting evaluation strategy. (A) First, a set of similar molecule SMILES is rendered into molecular images. These images are subsequently converted into fingerprints using each of the evaluated methods. (B) Second, a query SMILES string is transformed into a fingerprint. Its similarity is computed against all fingerprints within the dataset. Finally, based on the ranking of the similarity scores, the position of the correct molecule is reported.
@@ 7.3.1 Datasets and Metrics
We evaluate visual fingerprinting methods using five benchmark datasets: adenosine, camphor, cholesterol, limonene, and pyridine. The construction of these benchmarks is illustrated in Figure 5.
First, for each reference molecule (adenosine, camphor, cholesterol, limonene, and pyridine), we sample 500 SMILES from PubChem that have at least 90% structural similarity to the corresponding reference molecule. Second, we convert this sets of similar molecules to a sets of images by rendering them using RDKit. The resulting images undergo substantial data augmentation, including shifting, scaling, rotation, downscaling, and grid distortion. Third, each visual fingerprinting method converts these augmented images into molecular fingerprints. The methods evaluated are SubGrapher, OSRA with RDKit Daylight [19] or MHFP [22], and MolScribe with RDKit Daylight or MHFP. Then, to assess retrieval performance, we measure each method’s ability to correctly retrieve a molecular image using its fingerprint from a provided SMILES query. First, the query SMILES is converted into a fingerprint based on the evaluated method. Second, the similarity between this query fingerprint and all fingerprints in the dataset is computed. Third, the results are ranked based on similarity scores, and the position of the correct ground-truth molecule is recorded. Each benchmark is tested using 50 SMILES queries, and the average rank across these queries is reported. Intuitively, this setup evaluates the ability to retrieve a molecular image from a document based on a user-provided SMILES query under extreme conditions, as all molecular images in the dataset are highly similar and augmented. Although the test sets contain only 500 images, this design aims to simulate searches in larger collections.
@@ 7.3.2 State-of-the-art Comparison
Models Adenosine (500) Camphor (500) Cholesterol (500) Limonene (500) Pyridine (500)
@@
Table 2 compares visual fingerprinting methods for similarity search. On average, SubGrapher ranks first for camphor, cholesterol, and pyridine, and second for adenosine and limonene. Additionally, SubGrapher retrieves the correct reference molecule at an average rank of 95, significantly outperforming other methods. Given the challenging nature of the evaluation, we expect that, in practical scenarios, the target molecule would typically appear within the top 100 results, making manual inspection feasible. SubGrapher also performs more consistently across different types of reference molecules, suggesting that it handles structural variations reliably. Our results show that converting molecular images directly into fingerprints in a single step produces more distinctive fingerprints.
@@ 8 Conclusion
We present SubGrapher a learning-based approach for converting 2D chemical structure depictions to molecular fingerprints. SubGrapher detects functional groups and carbon backbones, assembling them into a graph representation, which is subsequently converted into a fingerprint. This fingerprint enables efficient substructure and molecule retrieval from collections of molecular images. Trained on synthetic images, SubGrapher demonstrates strong generalization capabilities, making it useful for large-scale molecular search. Our results highlight the benefits of strong mask supervision in molecular image recognition and demonstrate that directly converting images to fingerprints in a single step significantly improves molecular retrieval performance.

### v2

@@ 7.1.2 State-of-the-art Comparison
Figure 4: Substructure detection qualitative evaluation. Examples of predicted functional groups are shown for images from patent documents (JPO, USPTO-10K-L, USPTO-Markush) and a scientific journal (JOC).
@@
On the USPTO-Markush dataset, SubGrapher again outperforms other methods in substructure F1-score. This superior performance results from the limitations of competing models: OSRA and MolGrapher are trained exclusively on molecular images, whereas MolScribe and DECIMER handle Markush structures but support only limited Markush features, specifically variable groups represented as abbreviations. SubGrapher effectively manages complex Markush structures by focusing on relevant regions of the image and excluding Markush-related annotations.
@@ 7.2 Qualitative Evaluation
Figure 4 showcases examples of predicted molecules for images from various benchmark datasets. SubGrapher accurately recognizes functional groups in molecule images that contain captions or are of lower quality (Figure 4, row 1). It also recognizes functional groups in complex Markush structures (Figure 4, row 2) or unconventional drawings displayed in scientific publications (Figure 4, row 3). Additionally, our method maintain strong performances on extremely large molecules (Figure 4, row 4). Unlike image captioning methods such as DECIMER, our predictions preserve the spatial arrangement of substructures from the input image. This conveys useful information for human interpretation. Additional examples of predictions and typical failure cases are shown in the Supplementary Figure 2.
@@ 7.3 Visual Fingerprinting Evaluation
In this section, we evaluate SubGrapher for the retrieval of molecules or Markush structures in a collection of images.
Figure 5: Visual fingerprinting evaluation strategy. (A) First, a set of similar molecules are rendered into images. These images are subsequently converted into fingerprints using each of the evaluated methods. (B) Second, a molecule is converted into a fingerprint based on its SMILES. Its similarity is calculated against all fingerprints within the dataset. Finally, the correct molecule’s position is determined from the ranking of similarity scores.
@@ 7.3.1 Datasets and Metrics
We evaluate visual fingerprinting methods using five benchmarks of molecules: adenosine, camphor, cholesterol, limonene, and pyridine. The construction of these benchmarks is illustrated in Figure 5.
First, for each reference molecule (adenosine, camphor, cholesterol, limonene, and pyridine), we sample 500 SMILES from PubChem that have at least 90% structural similarity to the corresponding reference molecule. Here, the structural similarity is the Tanimoto [43] similarity computed between the PubChem fingerprints. Second, we convert this sets of similar molecules to a sets of images by rendering them using RDKit. The resulting images undergo substantial data augmentation, including scaling, rotation, downscaling, and grid distortion. Third, each visual fingerprinting method converts these augmented images into molecular fingerprints. The methods evaluated are SubGrapher, OSRA with RDKit Daylight [33] or MHFP [23], and MolScribe with RDKit Daylight or MHFP. Then, to assess retrieval performance, we measure each method’s ability to correctly retrieve a molecular image from a provided SMILES query. First, the query SMILES is converted into a fingerprint by following the same method than the evaluated fingerprint. Second, the similarity between this query fingerprint and all fingerprints in the dataset is computed. Third, the results are ranked based on similarity scores, and the position of the correct ground-truth molecule is recorded. RDKit fingerprint similarity is computed using the Tanimoto similarity on binary fingerprints, the MHFP similarity is computed using the Jaccard similarity, and the SVMF similarity is computed using the Euclidean distance. Each benchmark is tested using 50 SMILES queries, and the average rank across these queries is reported. This setup aims to evaluate the ability to retrieve depictions of a molecule in a large collection of documents based on a provided SMILES query. As all molecular images in the dataset are highly similar and augmented, despite the relatively small size of test sets, we aim to simulate searches in large collections. Example images from the benchmarks are shown in the Supplementary Figure 3.
@@ 7.3.2 State-of-the-art Comparison
Models
Adenosine
Camphor
Cholesterol
Limonene
Pyridine
@@
Table 2 compares visual fingerprinting methods for molecular retrieval. SubGrapher ranks first for datasets derived from camphor, cholesterol, and pyridine, and second for adenosine and limonene. In average over all benchmarks, SubGrapher retrieves the correct reference molecule at rank 95, significantly outperforming other methods. Given the challenging nature of the evaluation, we expect that, for a search in a large document collection, the target molecule image should appear at worse within the first 100 results. This reasonable number makes manual inspection feasible for critical search applications such as freedom-to-operate or prior-art search [44]. One reason for SubGrapher’s superior performance lies in how it handles uncertain predictions. For OCSR-based methods, all images converted to invalid SMILES are mapped to the same fingerprint. These identical and uninformative fingerprints degrade the ranking quality. In contrast, SubGrapher’s one-stage approach generates a distinctive fingerprint for each prediction, even when the prediction is uncertain. Overall, SubGrapher performs more consistently across different reference molecules than other methods, though its performance is slightly weaker on adenosine, cholesterol, and pyridine. For adenosine and cholesterol, this may be due to their larger size compared to the other compounds. For pyridine, the lower performance may come from the presence of charged molecules with multiple fragments in the dataset. SubGrapher currently does not recognize any substructures in single-atom fragments, leading to a loss of information. Besides, in the Supplementary Figure 4, we show that SubGrapher can distinguish positional isomers and homologous compounds, but not enantiomers. SubGrapher currently has the limitation of ignoring stereo-chemistry information. The Supplementary Table 3 also evaluate the impact of the augmentation level of the benchmarks on the SubGrapher performances. Overall, our results show that converting molecular images directly into fingerprints in a single step produces more distinctive fingerprints.
7.3.3 Qualitative Evaluation
Here, we demonstrate a use case of SubGrapher for retrieving a Markush structure within a patent document, as illustrated in Figure 6.
First, we extract molecule and Markush structure images from the patent ‘US20100016341A1’ using the DECIMER-Segmentation model [45], a deep neural network designed to segment chemical images from page images. Applying this model to the 54 pages of the document yields 356 extracted images. Next, each image is converted to its SVMF using SubGrapher. We then randomly select one of the extracted Markush structures and obtain its CXSMILES [34] using MarvinJS [46]. The CXSMILES is converted into its associated SVMF using ground-truth matched substructures.
Finally, we compare the query fingerprint against all SVMF in the dataset and rank the results by similarity. Figure 6 presents the top two matches along with their corresponding pages from the document. The ground-truth Markush structure image is correctly retrieved as the first match. This type of search can be particularly valuable for searching information in patent documents that goes beyond standard molecule depictions.
Figure 6: Visual fingerprinting qualitative evaluation. First, images are extracted from a patent document using a molecule image segmentation model. Then, these images are converted to SVMF using SubGrapher. Next, we select a query Markush structure and obtain the fingerprint associated to its structure. Finally, this query fingerprint is compared to all visual fingerprints of the dataset to obtain a ranking.
@@ 8 Conclusion
We present SubGrapher, a learning-based approach for converting 2D depictions of molecules and Markush structures into fingerprints. SubGrapher detects functional groups and carbon backbones, assembles them into a graph representation, and then converts this graph into a fingerprint. This fingerprint enables efficient substructure search and retrieval from collections of molecule and Markush structure images. Although trained solely on synthetic images, SubGrapher demonstrates strong generalization to real-world data. Our results highlight the advantages of pixel-level mask supervision for chemical image recognition. Unlike most existing methods, which first convert images to SMILES strings and then derive fingerprints from those SMILES, SubGrapher performs this conversion in a single step. While SMILES extraction remains essential for many applications, we show that a single-step approach achieves superior performance for retrieving molecules from image collections. SubGrapher also represents a step towards the retrieval of Markush structures from documents.
Declarations
8.1 Availability of data and materials
The SubGrapher code is available on GitHub: https://github.com/DS4SD/SubGrapher/.
The SubGrapher model weights are available on HuggingFace: https://huggingface.co/ds4sd/SubGrapher.
The visual fingerprinting benchmarks are available on HuggingFace: https://huggingface.co/datasets/ds4sd/SubGrapher-Datasets.
8.2 Authors’ contributions
L.M., G.I.M. and V.W conceptualized the substructure recognition model and fingerprinting method. L.M implemented the code. P.W.J.S and L.V.G supervised the work. L.M. wrote a draft of the manuscript. All authors revised and commented on the manuscript.
Supplementary Information:
SubGrapher: Visual Fingerprinting of Chemical Structures
Figure 7: Functional groups diversity. t-SNE visualization of 25 functional group clusters. Each cluster is represented by colored points and overlaid with the centroid molecule.
8.3 Supplementary Note 1
To assess the coverage of our substructures, we analyzed their presence in 122M molecules from PubChem. First, we evaluated how well our functional groups cover the heteroatoms in molecules that contain at least one heteroatom. We found that 97% of such molecules include at least one of our functional groups. On average, 77% of their heteroatoms are covered by at least one detected functional group. Second, we performed the same analysis for carbon backbones in molecules composed solely of carbon atoms. We found that 95% of these carbon-only molecules include at least one of our carbon patterns. On average, 71% of their atoms are captured by at least one carbon backbone from our list.
Expansion value Adenosine Camphor Cholesterol Limonene Pyridine
1d-1*d 125 80 126 89 112
0d0*d 111 74 106 82 105
0.05d0.05*d 111 74 106 81 103
0.1d0.1*d 110 73 106 81 103
0.25d0.25*d 110 73 106 82 104
Table 3: Expansion value sensitivity analysis. We compare the retrieval performance of SubGrapher for various expansion values of the detected bounding boxes on image datasets generated from adenosine, camphor, cholesterol, limonene, and pyridine. We report the average rank at which a query molecule is retrieved across 50 queries. $d$ denotes the diagonal length of the smallest detected box in the image.
Fingerprint Type Dimension Adenosine Camphor Cholesterol Limonene Pyridine
RDKit [33] Binary $1\times 4096$ 1465 392 696 230 249
HMFP [23] Integer $1\times 2048$ 2048 2048 2048 2048 2048
SVFP (Ours) Float
$1561\times 1561=$
$1\times 2436721$
43 20 35 15 7
Table 4: Fingerprints characteristics comparison. Comparison of fingerprints types, dimensions, and average number of non-zero coefficients computed on image retrieval benchmarks generated from adenosine, camphor, cholesterol, limonene, and pyridine. For RDKit and HMFP, the fingerprints are obtained using OSRA predictions.
Figure 8: Failure cases. Example of failure cases of SubGrapher on real-world data from JPO, USPTO-Markush and USPTO-10K-L. Typical failure cases include images containing captions (input 1), variable groups (input 2), long bonds (input 3), dashed bonds indicating stereo information (input 4), low resolution images (input 5) and inversions of similar substructures (input 6).
Figure 9: Benchmarks example images. Example images randomly selected from the benchmark sets generated from adenosine, camphor, cholesterol, limonene, and pyridine.
Figure 10: SVMF fingerprint discriminative capacity. SubGrapher’s predictions and SVMF fingerprints for example images representing (1) homologous compounds, (2) positional isomers, and (3) enantiomers. Cells in the SVMF fingerprints that differ between compounds are highlighted in red. SubGrapher distinguishes homologous compounds and positional isomers, but not enantiomers, since the substructures it detects lack stereochemistry information.
Augmentation level Adenosine Camphor Cholesterol Limonene Pyridine
Level 1 52 50 76 60 86
Level 2 110 73 106 81 103
Level 3 159 86 157 85 171
Table 5: Benchmarks augmentation analysis. We compare the retrieval performance of SubGrapher on different variants of the image datasets generated from adenosine, camphor, cholesterol, limonene, and pyridine. Each variant uses different augmentations levels as described in the Supplementary Note 2. We report the average rank at which a query molecule is retrieved across 50 queries.
8.4 Supplementary Note 2
To analyze the molecule retrieval results presented in the main manuscript, we perform evaluations on multiple augmented versions of the benchmarks. Table 5 compares the retrieval performance of SubGrapher on different variants of the image datasets generated from adenosine, camphor, cholesterol, limonene, and pyridine. Each benchmark is augmented using different levels of augmentations. Level 1 corresponds to applying the augmentations:
    Rotation with factor drawn between -0.1 and 0.1 (applied with a probability of 90%),
    Scaling with factor drawn between -0.4 and -0.3 (90%),
    Downscaling with a factor drawn between 0.5 and 0.8 (70%),
    Grid distortion with factor drawn between -0.1 and 0.1 (50%).
Level 2 is used in the main manuscript and corresponds to applying the augmentations:
    Rotation with factor drawn between -0.1 and 0.1 (90%),
    Scaling with factor drawn between -0.7 and -0.5 (90%),
    Downscaling with a factor drawn between 0.7 and 0.99 (70%),
    Grid distortion with factor drawn between -0.15 and 0.15 (50%).
Level 3 corresponds to applying the augmentations:
    Rotation with factor drawn between -0.15 and 0.15 (90%),
    Scaling with factor drawn between -0.8 and -0.6 (90%),
    Downscaling with a factor drawn between 0.8 and 0.99 (70%),
    Grid distortion with factor drawn between -0.2 and 0.2 (50%).
We observe that augmentation, particularly strong downscaling and grid distortion, has a significant impact on performance. This effect is especially pronounced for the benchmarks derived from adenosine, cholesterol, and pyridine. An explanation is that these benchmarks contain more heteroatoms than the limonene and camphor benchmarks, making them more sensitive to the loss of detail caused by downscaling.